OpenStudy (elisaneedshelp):

#12 REVIEW )))))))) @ganeshie8 can you help me again?

ganeshie8 (ganeshie8):

yes

OpenStudy (elisaneedshelp):

how do i solve this?

ganeshie8 (ganeshie8):

easy, use below : `altitude = "geometric mean" of segments of hypotenuse`

ganeshie8 (ganeshie8):

\(h = \sqrt{10 \times 25}\) simplify

OpenStudy (elisaneedshelp):

i got 15.8113883

ganeshie8 (ganeshie8):

Or u may use a proportion as all the triangles shown are similar : \[\dfrac{h}{10} = \dfrac{25}{h}\]

ganeshie8 (ganeshie8):

\(h = \sqrt{10\times 25} = 5\sqrt{10} \approx 15.81\)

OpenStudy (elisaneedshelp):

what about this 1? is there anything to solve?

ganeshie8 (ganeshie8):

u familiar wid this property : `tangent and radius intersect at 90 degrees` ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):